Solve the system of equations. 4x-y=6 2x+3y=-4
Answer by Mathtut(3670)

4x-y=6 ..........eq 1 rewrite to y=4x-6
2x+3y=-4.........eq 2
:
take y's value from eq 1 of 4x-6 and plug it into eq 2
:
2x+3(4x-6)=-4
:
2x+12x-18=-4
:
14x=14
:
x=1
:
now take x's found value and plug it back into either equation and solve for y
:
I will use rewrite of eq 1
:
y=4(1)-6=-2
:
(1,-2) is the solution